# AffectGPT-R1: Leveraging Reinforcement Learning for Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ 3.3.3 Reward Hacking
During reinforcement learning, we observe that the model tends to generate lengthy outputs $o^{a}$ containing numerous emotional words. The primary reason for this is that, in the calculation of EW-based metrics, synonyms are automatically disregarded (see Appendix A). Consequently, synonyms do not influence the metric scores, whereas predicting more words may instead increase the likelihood of matching the ground truth. This results in lengthy outputs, which in turn increases inference time and reduces the readability of $o^{a}$. This phenomenon, also referred to as reward hacking [everitt2017reinforcement], occurs when a model exploits loopholes or flaws in the reward function design to achieve high rewards while generating undesired behavior. To generate human-preferred outputs, we introduce length penalties based on the predicted label count, encouraging the model to maximize rewards with fewer prediction words. Specifically, we propose three types of penalties. Taking accuracy reward as an example, these penalties are defined as follows:
@@
Figure 1 visualizes the values of these penalties (see the reward hacking part). From $P_{1}(\cdot)$ to $P_{3}(\cdot)$, the penalty on prediction lengths becomes looser. In Section 5.3, we conduct experiments to evaluate the effects of different penalty strategies on emotion recognition performance and their efficacy in mitigating redundant predictions.
@@ 4.1 Corpus Description
AffectGPT-R1 contains two training phases: cold-start training and reinforcement learning. In the first phase, we employ a large-scale dataset to equip the model with emotion understanding and format alignment capabilities. For this purpose, we select MER-Caption+ [lian2025affectgpt], which contains 31K automatically annotated emotion descriptions. In the second phase, we need a fine-grained dataset to better address OV-MER. Thus, we choose MER2025-OV [lian2025mer], which includes 1,200 samples with human-annotated open-vocabulary emotion labels. For performance comparison, we evaluate the model on OV-MERD+ [lian2025affectgpt] and MER-UniBench [lian2025affectgpt], assessing its capabilities in open-vocabulary and generalized emotion understanding. Appendix D provides additional details about MER-UniBench. Since 200 samples overlap between MER2025-OV and OV-MERD+, we exclude these samples from MER2025-OV to ensure a fair comparison, resulting in 1,000 samples used in reinforcement learning. Table 1 presents statistics for these datasets. As shown, during cold-start training, we leverage both descriptive and open-vocabulary emotions provided by the dataset. However, in reinforcement learning, we exclusively use open-vocabulary emotions. This decision stems from the fact that the reward calculation process does not rely on human-annotated thinking results (see Section 3.3.2). Therefore, our dataset selection is well-suited to the distinct requirements of each training stage.
Table 1: Dataset statistics. We summarize the datasets used for training and testing. In this table, “Sen.”, “Bas.”, “OV”, and “Des.” are abbreviations for sentiment labels, basic emotions, open-vocabulary emotions, and descriptive emotions, respectively.
@@
Cold-start training
MER-Caption+ [lian2025affectgpt] 31,327 $\times$ $\times$ $\surd$ $\surd$
Reinforcement learning
MER2025-OV [lian2025mer] 1,000 $\times$ $\times$ $\surd$ $\times$
@@
OV-MERD+ [lian2025affectgpt] 532 $\times$ $\times$ $\surd$ $\times$
MER-UniBench [lian2025affectgpt] 12,799 $\surd$ $\surd$ $\surd$ $\times$
@@ 4.2 Implementation Details
During training, we set the learning rate to 1e-5 and the maximum number of epochs to 60. The entire implementation is conducted using PyTorch, and the code is executed on an A100 GPU with 80 GB of memory. Due to memory constraints, we use a batch size of 3 for cold-start training and a batch size of 1 for reinforcement learning. For the model architecture, we adopt the default hyperparameters of AffectGPT [lian2025affectgpt]. This setup ensures a fair comparison between AffectGPT and AffectGPT-R1, better highlighting the effectiveness of our proposed framework.
@@ 5 Results and Discussion
In this section, we first analyze the effectiveness of different reward functions, the impact of the output mode, and the effect of varying length penalties. Next, we compare the performance of AffectGPT-R1 with other baselines on MER-UniBench, demonstrating the effectiveness of our method for generalized emotion understanding. Finally, we conduct a series of ablation studies to investigate the influence of the choice of $G$, as well as the amount of data used in cold-start training and reinforcement learning.
@@ 5.1 Reward Selection
In the single-reward scenarios, we observe that the accuracy reward and dual reward are the most effective, significantly improving performance compared to the model without reinforcement learning. In contrast, other rewards lead to performance degradation, indicating potential reward hacking. The superior performance of the accuracy reward and dual reward stems from their direct optimization of EW-based metrics: one on $o^{a}$ (answer) and one on $o^{t}$ (thinking). This highlights the importance of directly aligning reinforcement learning with evaluation metrics in OV-MER. Meanwhile, the format reward and alignment reward show limited impact, because these aspects are already well-learned during the cold-start phase. The weak performance of the perception reward may be due to current MLLMs’ inability to accurately decode human preferences, which could introduce reward noise [lian2025emoprefer]. Interestingly, the accuracy reward and dual reward yield similar performance, suggesting that while EW-based metrics are crucial, the choice of optimizing on thinking $o^{t}$ or answer $o^{a}$ has minimal impact. Therefore, in the next experiment, we use the accuracy reward by default.
@@
accuracy format alignment dual preception
@@ 5.2 Output Mode Analysis
In single-reward setups (see Table 2), the most effective rewards are the accuracy reward and dual reward, both of which are directly optimized on EW-based metrics. This observation leads to a heuristic hypothesis: Can we modify the output mode to generate and optimize solely on $o^{a}$, thereby making the model more focused on EW-based metrics? Specifically, for cold-start training, we remove the thinking process and require the output to contain only $o^{a}$. For reinforcement learning, we optimize the accuracy reward on $o^{a}$ directly. We denote this modified model as AffectGPT-R1 (w/o think) and the original model in Figure 1 as AffectGPT-R1 (w/ think). Experimental results in Table 3 show that AffectGPT-R1 (w/o think) achieves better performance. This improvement can be attributed to two reasons. Firstly, our thinking process is derived from coarse-grained datasets, where dataset noise may introduce incorrect reasoning and degrade performance. Secondly, including additional thinking content diverts the model’s attention from addressing OV-MER, thereby increasing training difficulty.
@@ 5.3 Impact of Different Penalties
During inference, the model tends to generate lengthy responses with synonyms. This behavior primarily stems from the suboptimal design of the reward function. To mitigate this issue, we introduce three types of penalties in Section 3.3.3, where the penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$. In this section, we evaluate the effectiveness of these penalties across three dimensions: (1) their impact on performance, (2) their ability to prevent overlong outputs, and (3) whether they introduce reward degradation. As shown in Table 4, all three penalties successfully address overlong outputs without compromising performance. However, the strict penalty, $P_{1}(\cdot)$, leads to reward degradation. Therefore, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are better choices.
Table 4: Impact of different penalties. The penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$. Here, “overlong” indicates the presence of overlong outputs, and “degradation” indicates the presence of reward degradation. We seek a length penalty that preserves accuracy, prevents overlong outputs, and avoids causing reward degradation. Therefore, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are better choices.
Penalty Accuracy Overlong Degradation
@@
5.4 Main Results on MER-UniBench
The above sections primarily focus on open-vocabulary emotions. In this section, we broaden the evaluation scope to MER-UniBench [lian2025affectgpt], a unified benchmark encompassing three typical tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection. Experimental results are presented in Table 5. For fairness, all comparisons are conducted under the same input modality. These results demonstrate that AffectGPT-R1 achieves notable performance improvements in basic and fine-grained emotion recognition but exhibits slight degradation in sentiment analysis. This discrepancy arises because fine-grained and basic MER primarily focus on discrete emotion words (e.g., happy), whereas sentiment analysis centers on emotion polarity (e.g., positive and negative). Consequently, fine-grained and basic MER are more closely related to each other. Based on the average results across three tasks, AffectGPT-R1 achieves state-of-the-art performance on MER-UniBench, demonstrating the advantages of reinforcement learning for multimodal emotion understanding.
@@ 5.4 Main Results on MER-UniBench
OneLLM [han2024onellm] 64.01 54.09 63.39 61.98 25.52 17.21 28.32 33.44 22.25 41.14
SECap [xu2024secap] 55.76 54.18 59.51 57.41 40.95 52.46 25.56 36.92 36.97 46.64
PandaGPT [su2023pandagpt] 66.06 61.33 62.93 58.88 33.57 39.04 31.91 36.55 31.33 46.84
Qwen-Audio [chu2023qwen] 70.09 46.90 70.73 65.26 41.85 31.61 49.09 35.47 32.36 49.26
SALMONN [tang2023salmonn] 81.00 67.03 68.69 65.93 55.53 45.38 45.62 46.84 45.00 57.89
AffectGPT [lian2025affectgpt] 83.46 80.74 82.99 83.75 72.94 73.41 56.63 55.68 59.98 72.18
AffectGPT-R1 80.13 80.01 84.49 82.31 81.69 93.49 63.74 63.85 65.49 77.24
@@
Otter [li2025otter] 52.89 50.44 57.56 53.12 16.41 14.65 22.57 29.08 16.63 34.82
Video-LLaVA [lin2024video] 56.37 61.64 53.28 57.45 36.93 30.25 30.73 38.95 34.00 44.40
PandaGPT [su2023pandagpt] 58.50 64.25 62.07 65.25 39.13 47.16 38.33 47.21 35.07 50.77
Video-ChatGPT [maaz2024video] 54.42 63.12 64.82 65.80 44.86 46.80 37.33 56.83 39.80 52.64
VideoChat2 [li2024mvbench] 66.84 54.32 69.49 70.66 33.67 54.50 36.64 48.70 39.21 52.67
LLaMA-VID [li2024llama] 61.78 63.89 69.35 67.48 50.72 57.60 42.75 46.02 45.01 56.07
VideoChat [li2023videochat] 65.13 63.61 69.52 72.14 48.73 57.30 41.11 48.38 44.52 56.71
Chat-UniVi [jin2024chat] 54.53 63.18 68.15 66.36 57.62 65.67 45.61 52.37 48.00 57.94
mPLUG-Owl [ye2023mplug] 72.40 72.91 72.13 75.00 56.86 59.89 49.11 55.54 48.18 62.45
AffectGPT [lian2025affectgpt] 82.39 81.57 87.20 86.29 74.58 75.29 57.63 62.19 61.65 74.31
AffectGPT-R1 78.78 79.07 85.91 85.85 77.72 85.29 61.09 67.42 62.42 75.95
@@
PandaGPT [su2023pandagpt] 61.92 67.61 68.38 67.23 40.21 51.89 37.88 44.04 37.12 52.92
R1-Omni [zhao2025r1] 58.02 56.48 71.82 68.58 64.17 67.43 43.20 51.58 55.24 59.61
Emotion-LLaMA [cheng2024emotion] 66.13 67.66 78.32 77.23 59.38 73.62 46.76 55.47 52.97 64.17
AffectGPT [lian2025affectgpt] 81.30 80.90 88.49 86.18 78.54 78.80 55.65 60.54 62.52 74.77
AffectGPT-R1 79.65 80.18 87.26 85.75 84.51 93.13 66.71 74.26 68.39 79.98
@@ 5.5 Ablation Study
Figure 2: Left: Impact of $G$. “–” denotes the baseline model without reinforcement learning. Mid: Impact of data amount in cold-start training. This figure presents the performance of models trained with varying amounts of cold-start data and analyzes their influence on subsequent reinforcement learning. Right: Impact of data amount in reinforcement learning. “–” denotes the baseline model without reinforcement learning. This figure shows that using a larger amount of data in reinforcement learning leads to better performance.
@@
GRPO requires sampling $G$ responses for group normalization (see Section 3.3.1). Figure 2 (left) reveals the impact of $G$. Due to GPU memory constraints, we limit our comparison to values of $G$ ranging from 2 to 16 and do not evaluate larger values. Experimental results demonstrate that, regardless of the chosen $G$, the model with reinforcement learning outperforms the one without reinforcement learning, demonstrating the effectiveness of the R1-style strategy in OV-MER. However, the selection of $G$ also matters. A small $G$ (e.g., $G=2$) fails to fully leverage the advantages of GRPO, primarily because insufficient sampling limits the likelihood of obtaining optimal results. Conversely, a larger $G$ (e.g., $G\geq 4$) ensures the effectiveness of GRPO, but further increasing $G$ (e.g., $G=16$) yields no significant improvement. Given that a larger $G$ also imposes higher GPU memory requirements, we adopt $G=4$ as our default setting.
Data Amount in Cold-start Training.
AffectGPT-R1 consists of two stages: cold-start training (denoted as $s_{1}$) and reinforcement learning (denoted as $s_{2}$). In this section, we investigate the impact of the amount of cold-start data on $s_{1}$ and $s_{2}$. Experimental results in Figure 2 (mid) demonstrate that increasing the amount of cold-start data improves performance in $s_{1}$. Meanwhile, we observe that the performance after reinforcement learning is closely tied to the performance of the base model from $s_{1}$, and more cold-start data also enhances $s_{2}$ performance. The main reason lies in that reinforcement learning predominantly serves to reinforce correct answers rather than infuse novel knowledge [liu2025understanding]. Therefore, a high-quality base model is essential for maximizing the effectiveness of reinforcement learning.
Data Amount in Reinforcement Learning.
This section investigates the influence of data quantity on reinforcement training. Specifically, we start from the same model obtained through cold-start training (i.e., the model trained on 100% data in Figure 2 (mid)) and evaluate its performance with different amounts of reinforcement learning data. Experimental results in Figure 2 (right) demonstrate that even using just 10 samples for reinforcement learning leads to performance improvements, and further increasing the data amount yields more substantial gains. These phenomenon indicates that AffectGPT-R1 has not yet reached its upper-bound performance. In the future, we plan to collect more data to further improve the performance of AffectGPT-R1.

### v3

@@ 3.3.3 Reward Hacking
During reinforcement learning, we observe that the model tends to generate lengthy outputs $o^{a}$ containing numerous emotional words. The primary reason for this is that, in the calculation of EW-based metrics, synonyms are automatically disregarded (see Appendix D). Consequently, synonyms do not influence the metric scores, whereas predicting more words may instead increase the likelihood of matching the ground truth. This results in lengthy outputs, which in turn increases inference time and reduces the readability of $o^{a}$. This phenomenon, also referred to as reward hacking (Everitt et al., 2017), occurs when a model exploits loopholes or flaws in the reward function design to achieve high rewards while generating undesired behavior. To generate human-preferred outputs, we introduce length penalties based on the predicted label count, encouraging the model to maximize rewards with fewer prediction words. Specifically, we propose three types of penalties. Taking accuracy reward as an example, these penalties are defined as follows:
@@
Figure 1 visualizes the values of these penalties (see the reward hacking part). From $P_{1}(\cdot)$ to $P_{3}(\cdot)$, the penalty on prediction lengths becomes looser. In Section 5.3, we conduct experiments to evaluate the effects of different penalty strategies on recognition performance and their efficacy in mitigating redundant predictions.
@@ 4.1 Corpus Description
AffectGPT-R1 consists of two training phases: cold-start training and reinforcement learning. In the first phase, we utilize a large-scale dataset to endow the model with emotion understanding and format alignment capabilities. For this purpose, we select MER-Caption+ (Lian et al., 2025a), which contains 31K samples annotated with rich emotion descriptions and open-vocabulary labels. In the second phase, we require a high-quality dataset specifically designed for OV-MER. Accordingly, we choose MER2025-OV (Lian et al., 2025b), which features human-annotated, high-quality open-vocabulary labels. For evaluation, we assess the model’s performance on OV-MERD+ (Lian et al., 2025a) and MER-UniBench (Lian et al., 2025a), measuring its capabilities in open-vocabulary and generalized emotion understanding. We conduct a rigorous overlap check to ensure there are no overlapping samples between the training and testing data. More details are provided in Appendix G.
Table 1 presents statistics for these datasets. As shown, during cold-start training, we leverage both descriptive and open-vocabulary emotions provided by the dataset. However, in reinforcement learning, we exclusively use open-vocabulary emotions. This decision stems from the fact that the reward calculation process does not rely on human-annotated thinking results (see Section 3.3.2). Therefore, our dataset selection is well-suited to the distinct requirements of each training stage.
Table 1: Dataset statistics. We summarize the datasets used for training and testing. In this table, “Sen.”, “Bas.”, “OV”, and “Des.” are abbreviations for sentiment labels, basic emotions, open-vocabulary emotions, and descriptive emotions, respectively. We rigorously verify that there is no overlap between the training and testing data. More details can be found in Appendix G.
@@
Training Phase 1: Cold-start training
MER-Caption+ (Lian et al., 2025a) 31,327 $\times$ $\times$ $\surd$ $\surd$
Training Phase 2: Reinforcement learning
MER2025-OV (Lian et al., 2025b) 1,000 $\times$ $\times$ $\surd$ $\times$
@@
OV-MERD+ (Lian et al., 2025a) 532 $\times$ $\times$ $\surd$ $\times$
MER-UniBench (Lian et al., 2025a) 12,799 $\surd$ $\surd$ $\surd$ $\times$
@@ 4.2 Implementation Details
During training, we set the learning rate to 1e-5 and the maximum number of epochs to 60. The entire implementation is conducted using PyTorch, and the code is executed on an A100 GPU with 80 GB of memory. Due to memory constraints, we use a batch size of 3 for cold-start training and a batch size of 1 for reinforcement learning. For the model architecture, we adopt the default hyperparameters of AffectGPT (Lian et al., 2025a). This setup ensures a fair comparison between AffectGPT and AffectGPT-R1, better highlighting the effectiveness of our proposed framework.
@@ 5.1 Reward Selection
In the single-reward scenarios, we observe that the accuracy reward and dual reward are the most effective, significantly improving performance compared to the model without reinforcement learning. In contrast, other rewards lead to performance degradation, indicating potential reward hacking. The superior performance of the accuracy reward and dual reward stems from their direct optimization of EW-based metrics: one on $o^{a}$ (answer) and one on $o^{t}$ (thinking). This highlights the importance of directly aligning reinforcement learning with evaluation metrics in OV-MER. Meanwhile, the format reward and alignment reward show limited impact, because these aspects are already well-learned during the cold-start phase. The weak performance of the perception reward may be due to current MLLMs’ inability to accurately decode human preferences, which could introduce reward noise (Lian et al., 2026a). Interestingly, the accuracy reward and dual reward yield similar performance, suggesting that while EW-based metrics are crucial, the choice of optimizing on thinking $o^{t}$ or answer $o^{a}$ has minimal impact. Therefore, in the next experiment, we use the accuracy reward by default.
@@
accuracy format alignment dual perception
@@ 5.2 Output Mode Analysis
In single-reward setups (see Table 2), the most effective rewards are the accuracy reward and dual reward, both of which are directly optimized based on EW metrics. This observation motivates a heuristic hypothesis: Could we modify the output mode to generate and optimize solely on $o^{a}$, thereby encouraging the model to focus more on EW-based metrics? Specifically, during cold-start training, we eliminate the thinking process and require the output to contain only $o^{a}$. During reinforcement learning, we optimize the accuracy reward directly on $o^{a}$. We refer to the modified model as AffectGPT-R1 (w/o think), and the original model shown in Figure 1 as AffectGPT-R1 (w/ think). In Table 3, we observe that removing the thinking process yields improved performance. This improvement can be attributed to two factors. First, our thinking process is derived from coarse-grained datasets, where inherent noise may lead to flawed reasoning and consequently degrade performance. Second, incorporating additional thinking content distracts the model from solving OV-MER, thereby increasing training difficulty. Regardless of whether the thinking process proves useful, this work represents a pioneering work in applying reinforcement learning to OV-MER. Future work will involve designing high-quality cold-start data and conducting an in-depth analysis to determine which samples or emotional states require additional thinking processes.
@@ 5.3 Impact of Different Penalties
During inference, the model tends to generate lengthy predictions containing redundant emotion words. This behavior stems from the suboptimal design of the reward function. To mitigate this issue, we introduce three types of penalties in Section 3.3.3, where the penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$. In this section, we evaluate the effectiveness of these penalties across three dimensions: (1) their impact on performance, (2) their ability to prevent overlong outputs, and (3) whether they introduce reward degradation. As shown in Table 4, all three penalties successfully address overlong outputs without compromising performance. However, the strict penalty, $P_{1}(\cdot)$, leads to reward degradation. Therefore, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are better choices.
Table 4: Impact of different penalties. Our goal is to identify a penalty that maintains model performance, prevents overlong outputs, and avoids introducing reward degradation. Based on these criteria, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are more suitable choices.
Penalty Performance Overlong Degradation
@@
OneLLM (Han et al., 2024) 64.01 54.09 63.39 61.98 25.52 17.21 28.32 33.44 22.25 41.14
SECap (Xu et al., 2024) 55.76 54.18 59.51 57.41 40.95 52.46 25.56 36.92 36.97 46.64
PandaGPT (Su et al., 2023) 66.06 61.33 62.93 58.88 33.57 39.04 31.91 36.55 31.33 46.84
Qwen-Audio (Chu et al., 2023) 70.09 46.90 70.73 65.26 41.85 31.61 49.09 35.47 32.36 49.26
SALMONN (Tang et al., 2023) 81.00 67.03 68.69 65.93 55.53 45.38 45.62 46.84 45.00 57.89
AffectGPT (Lian et al., 2025a) 83.46 80.74 82.99 83.75 72.94 73.41 56.63 55.68 59.98 72.18
AffectGPT-R1 (Ours) 80.13 80.01 84.49 82.31 81.69 93.49 63.74 63.85 65.49 77.24
@@
Otter (Li et al., 2025a) 52.89 50.44 57.56 53.12 16.41 14.65 22.57 29.08 16.63 34.82
Video-LLaVA (Lin et al., 2024) 56.37 61.64 53.28 57.45 36.93 30.25 30.73 38.95 34.00 44.40
PandaGPT (Su et al., 2023) 58.50 64.25 62.07 65.25 39.13 47.16 38.33 47.21 35.07 50.77
Video-ChatGPT (Maaz et al., 2024) 54.42 63.12 64.82 65.80 44.86 46.80 37.33 56.83 39.80 52.64
VideoChat2 (Li et al., 2024a) 66.84 54.32 69.49 70.66 33.67 54.50 36.64 48.70 39.21 52.67
LLaMA-VID (Li et al., 2024b) 61.78 63.89 69.35 67.48 50.72 57.60 42.75 46.02 45.01 56.07
VideoChat (Li et al., 2025b) 65.13 63.61 69.52 72.14 48.73 57.30 41.11 48.38 44.52 56.71
Chat-UniVi (Jin et al., 2024) 54.53 63.18 68.15 66.36 57.62 65.67 45.61 52.37 48.00 57.94
mPLUG-Owl (Ye et al., 2023) 72.40 72.91 72.13 75.00 56.86 59.89 49.11 55.54 48.18 62.45
AffectGPT (Lian et al., 2025a) 82.39 81.57 87.20 86.29 74.58 75.29 57.63 62.19 61.65 74.31
AffectGPT-R1 (Ours) 78.78 79.07 85.91 85.85 77.72 85.29 61.09 67.42 62.42 75.95
@@
PandaGPT (Su et al., 2023) 61.92 67.61 68.38 67.23 40.21 51.89 37.88 44.04 37.12 52.92
R1-Omni (Zhao et al., 2025) 58.02 56.48 71.82 68.58 64.17 67.43 43.20 51.58 55.24 59.61
Emotion-LLaMA (Cheng et al., 2024) 66.13 67.66 78.32 77.23 59.38 73.62 46.76 55.47 52.97 64.17
AffectGPT (Lian et al., 2025a) 81.30 80.90 88.49 86.18 78.54 78.80 55.65 60.54 62.52 74.77
AffectGPT-R1 (Ours) 79.39 79.24 88.25 84.97 84.32 93.75 63.12 74.26 68.05 79.48
Figure 2: Left: Impact of data amount in cold-start training. This figure shows the performance of models trained with varying amounts of cold-start data and analyzes their influence on subsequent reinforcement learning. Mid: Impact of data amount in reinforcement learning. “–” denotes the baseline model without reinforcement learning. This figure shows that using a larger amount of data in reinforcement learning leads to better performance. Right: Impact of $G$. “–” denotes the baseline model without reinforcement learning.
5.4 Main Results on MER-UniBench
Global-level Analysis.
The preceding sections focus on open-vocabulary emotion recognition. In this section, we extend our evaluation to MER-UniBench (Lian et al., 2025a), a comprehensive benchmark that encompasses three tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection. Appendix G provides additional details about this benchmark. Experimental results are summarized in Table 5. To ensure a fair comparison, all methods are evaluated using the same input modality. Based on the average performance across three tasks, AffectGPT-R1 achieves state-of-the-art results on MER-UniBench, demonstrating the effectiveness of reinforcement learning in enhancing emotion understanding.
Task-level Analysis.
Through further analysis of each task, AffectGPT-R1 demonstrates significant performance improvements in both basic and fine-grained emotion recognition, but shows a slight decline in sentiment analysis. This discrepancy stems from the fact that this work takes OV-MER as its primary focus and specifically tailors reward functions for this task. OV-MER aims to recognize discrete emotion words (e.g., happy and surprise), which closely align with the objectives of both basic and fine-grained emotion recognition. However, it differs somewhat from sentiment analysis, which targets emotional polarity (e.g., positive and negative). This distinction results in the reinforcement learning-optimized model, AffectGPT-R1, performing slightly worse on sentiment analysis compared to the non-optimized AffectGPT. Future work involves incorporating sentiment-related rewards into policy updates.
@@ 5.5 Ablation Study
Data Amount in Cold-start Training.
AffectGPT-R1 consists of two stages: cold-start training (denoted as $s_{1}$) and reinforcement learning (denoted as $s_{2}$). In this section, we examine the impact of the amount of cold-start data on both $s_{1}$ and $s_{2}$. Experimental results shown in Figure 2 (left) indicate that increasing the amount of cold-start data improves performance in $s_{1}$. Furthermore, the performance after reinforcement learning strongly correlates with the performance of the base model from $s_{1}$, and additional cold-start data also benefits $s_{2}$. This is primarily because reinforcement learning mainly reinforces correct responses rather than introducing new knowledge (Liu et al., 2025). Therefore, a high-quality base model is crucial for maximizing the effectiveness of reinforcement learning.
Data Amount in Reinforcement Learning.
This section reveals the impact of data quantity on reinforcement learning. Specifically, starting from the same model obtained via cold-start training (i.e., the model trained on 100% data, as shown in the left part of Figure 2), we evaluate its performance using varying amounts of reinforcement learning data. Experimental results are shown in Figure 2 (mid). We observe that even using just 10 samples for reinforcement learning leads to performance improvements, confirming the effectiveness of reinforcement learning in addressing OV-MER. Moreover, increasing the amount of training data leads to greater performance gains. These phenomenon indicates that AffectGPT-R1 has not yet reached its upper-bound performance. In the future, we plan to collect more data to further improve the performance of AffectGPT-R1.
@@
GRPO requires sampling $G$ outputs for group normalization. Figure 2 (right) reveals the impact of $G$. Due to GPU memory limitations, we restrict our comparisons to values of $G$ between 2 and 16 and do not evaluate larger values. Experimental results show that, regardless of the specific value of $G$, the model trained with reinforcement learning consistently outperforms the one without it, confirming the effectiveness of our approach. Meanwhile, the selection of $G$ also matters. A small $G$ does not fully exploit the benefits of GRPO, mainly because limited sampling reduces the chances of identifying optimal outcomes. In contrast, a larger $G$ (e.g., $G\geq 4$) ensures the effectiveness of GRPO, but further increasing $G$ yields no significant improvement. Since a larger $G$ also increases computational cost, we set $G=4$ as the default configuration.